If four times a first number is decreased by three times a second number, the result is 0. The sum of the two numbers is -7.
Answer by ankor@dixie-net.com(22740)

Two numbers, a & b
:
If four times a first number is decreased by three times a second number, the result is 0.
4a - 3b = 0
:
The sum of the two numbers is -7.
a + b = -7
a = -b - 7
:
Replace a with (-b-7) in the first equation
4(-b-7) - 3b = 0
-4b - 28 - 3b = 0
-7b = + 28
-7b = +28
b = 28/-7
b = -4
:
Find a
a = -(-4) - 7
a = +4 - 7
a = -3
;
See if that works in "If four times a first number is decreased by three times a second number, the result is 0. "
4(-3) - 3(-4) =
-12 + 12 = 0